OpenStudy (anonymous):

how do i solve this:(x/x+3) + (2/5)=0

OpenStudy (anonymous):

remove the denominators first (x+3) and 5 has LCM as 5(x+3) multiply 5(x+3) to both sides. (x/x+3)5(x+3) + (2/5)5(x+3)=0 5x+2(x+3)=0 5x+2x+6=0 7x+6=0 7x+6-6=0-6 7x=-6 x=-6/7